OpenStudy (anonymous):

find the average value of the given function f(X) over the specified interval a<=x<=b f(x)= (e^x-e^-x)/(e^x+e^-x) over 0<=x<= ln 3

jimthompson5910 (jim_thompson5910):

You need to compute the integral of f(x) and multiply it by 1/(b-a) where 'a' and 'b' are the endpoints of the interval

jimthompson5910 (jim_thompson5910):

I recommend letting u = e^x + e^(-x) which means du = (e^x - e^(-x))dx

OpenStudy (loser66):